# Quantum property testing in sparse directed graphs

### v2

@@ 5.3 Other maximal hard-to-test problems
As we mentioned in the introduction, there are several other problems in the bounded degree graph model, which are maximally hard-to-test. Moreover, their lower bounds stem from similar ideas as the E$(3,c)$LIN-2 and $3$-colorability lower bounds, as mentioned in [yoshida2010query, goldreich2020testing]. Following the same path as in the previous subsection, we also obtain $\Omega(N)$ quantum query lower bounds for all these problems. For brevity, we only present the theorem statements below and omit their proofs.
@@ 6 Conclusion
Appendix A Deferred material from Section 4
@@ A.1 Dual polynomial for THR
We now restate Lemma 4.16 before proving it. See 4.16
@@
With the construction of $\psi$ described in Section A.1, Proposition A.2 (with $T=N$) ensures that Item 1 is satisfied. This way, we know that $\|\psi\|_{1}=1$, and that $\mathrm{phd}(\psi)\geq c_{1}\sqrt{k^{-1}N^{1-1/k}}$. From Proposition 4.15 we know that $\|\phi\|_{1}=1$ and $\mathrm{phd}(\phi)\geq 1$. Using item 1 of Proposition A.3 we obtain $\|\phi\star\psi\|_{1}=1$, and using Item 2 we get $\mathrm{phd}(\phi\star\psi)\geq c_{1}\sqrt{k^{-1}N^{1-1/k}}$.
From Proposition A.2 we know that the function $\omega$ that is used to define $\psi$ satisfies the $(\alpha,\beta)$-decay condition for some constant $\alpha=(2k)^{k}$ and $\beta=c_{2}/\sqrt{kN^{1+1/k}}$.
This way, we can use Proposition A.4 to obtain the function $\zeta$ we wanted for Item 2. Indeed, our functions $\psi$ and $\phi$ satisfy all the conditions of the lemma with pure high degree lower bounded by $\Delta=c_{1}\sqrt{k^{-1}N^{1-1/k}}$; and with our parameters $\alpha$, $\beta$ we obtain $\Delta^{\prime}=c_{2}(2k)^{k/2}R/(4\ln^{2}(R)\sqrt{kN^{1+1/k}})\in\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$. ∎
A.3 Proof of Proposition 4.17
@@
First part of Proposition 4.17
@@
We introduce new variables $z_{i}$ that will be compared to $\textnormal{sgn}(\psi(x_{i}))$. Using A.5, the probability of picking a $z\in\{-1,1\}^{R}$ from the uniform distribution such that $z$ corresponds to the vector of the signs is $\frac{1}{2^{R}}$. Thus previous term can be rewritten as
@@
Second part of Proposition 4.17
@@
Using A.5, we can first notice that for any $b\in\{-1,1\}$, the probability that an $x_{i}$ sampled from $\lambda$ is a false $b$ (i.e. false positive if $b=1$ and false negative if $b=-1$) is as follows, where by convention $D_{+1}=D_{+}$ and $D_{-1}=D_{-}$:
@@
Therefore, if $z_{i}=\textnormal{sgn}(\psi(x_{i}))$ and $x_{i}$ is a false $z_{i}$, it means that $z_{i}$ should be flipped to get $h(x_{i})$. Let $y_{i}\in\{-1,1\}$ denote whether we flip $z_{i}$. As $x_{i}$ is a false $z_{i}$ with probability $2\sum_{x_{i}\in D_{z_{i}}}|\psi(x_{i})|$, this is the probability with which we should flip $z_{i}$, i.e., the probability that $y_{i}=-1$.
@@
Appendix B Proofs from Section 5: hardness of testing E$(3,c)$LIN-2
In this section, we will present a sketch of the proof of the lower bound of testing E$(3,c)$LIN-2. Our main result is stated as follows.
@@ Appendix B Proofs from Section 5: hardness of testing E$(3,c)$LIN-2
Let us consider the matrix $A\in\{0,1\}^{cN\times N}$ as mentioned in Lemma 5.4. Based on the matrix $A$, the hard-to-distinguish distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ are as follows:
@@
Note that the system of linear equations in $D_{\mathrm{yes}}$ is satisfiable by setting $x=z$. On the other hand, for the system of linear equations corresponding to $D_{\mathrm{no}}$, vector $y$ is uniformly random. Thus, with high probability, vector $Az-y$ has large Hamming weight for any $z\in\{0,1\}^{N}$, and therefore the system of linear equations $Ax=y$ is far from being satisfiable. The formal proof is in [bogdanov2002lower, Lemma 18]. ∎

### v3

@@ 5.3 Other maximal hard-to-test problems
As we mentioned in the introduction, there are several other problems in the bounded degree graph model, which are maximally hard-to-test. Moreover, their lower bounds stem from similar ideas as the E(3,c)(3,c)LIN-2 and $3$-colorability lower bounds, as mentioned in [yoshida2010query, goldreich2020testing]. Following the same path as in the previous subsection, we also obtain $\Omega(N)$ quantum query lower bounds for all these problems. For brevity, we only present the theorem statements below and omit their proofs.
@@ 6 Conclusion
Appendix A Deferred material from Section˜4
@@ A.1 Dual polynomial for THR
We now restate Lemma˜4.16 before proving it. See 4.16
@@
With the construction of $\psi$ described in Section˜A.1, Proposition˜A.2 (with $T=N$) ensures that Item 1 is satisfied. This way, we know that $\|\psi\|_{1}=1$, and that $\mathrm{phd}(\psi)\geq c_{1}\sqrt{k^{-1}N^{1-1/k}}$. From Proposition˜4.15 we know that $\|\phi\|_{1}=1$ and $\mathrm{phd}(\phi)\geq 1$. Using item 1 of Proposition˜A.3 we obtain $\|\phi\star\psi\|_{1}=1$, and using Item 2 we get $\mathrm{phd}(\phi\star\psi)\geq c_{1}\sqrt{k^{-1}N^{1-1/k}}$.
From Proposition˜A.2 we know that the function $\omega$ that is used to define $\psi$ satisfies the $(\alpha,\beta)$-decay condition for some constant $\alpha=(2k)^{k}$ and $\beta=c_{2}/\sqrt{kN^{1+1/k}}$.
This way, we can use Proposition˜A.4 to obtain the function $\zeta$ we wanted for Item 2. Indeed, our functions $\psi$ and $\phi$ satisfy all the conditions of the lemma with pure high degree lower bounded by $\Delta=c_{1}\sqrt{k^{-1}N^{1-1/k}}$; and with our parameters $\alpha$, $\beta$ we obtain $\Delta^{\prime}=c_{2}(2k)^{k/2}R/(4\ln^{2}(R)\sqrt{kN^{1+1/k}})\in\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$. ∎
A.3 Proof of Proposition˜4.17
@@
First part of Proposition˜4.17
@@
We introduce new variables $z_{i}$ that will be compared to $\textnormal{sgn}(\psi(x_{i}))$. Using ˜A.5, the probability of picking a $z\in\{-1,1\}^{R}$ from the uniform distribution such that $z$ corresponds to the vector of the signs is $\frac{1}{2^{R}}$. Thus previous term can be rewritten as
@@
Second part of Proposition˜4.17
@@
Using ˜A.5, we can first notice that for any $b\in\{-1,1\}$, the probability that an $x_{i}$ sampled from $\lambda$ is a false $b$ (i.e. false positive if $b=1$ and false negative if b=1b=-1) is as follows, where by convention $D_{+1}=D_{+}$ and $D_{-1}=D_{-}$:
@@
Therefore, if $z_{i}=\textnormal{sgn}(\psi(x_{i}))$ and $x_{i}$ is a false $z_{i}$, it means that $z_{i}$ should be flipped to get $h(x_{i})$. Let $y_{i}\in\{-1,1\}$ denote whether we flip $z_{i}$. As $x_{i}$ is a false $z_{i}$ with probability $2\sum_{x_{i}\in D_{z_{i}}}|\psi(x_{i})|$, this is the probability with which we should flip $z_{i}$, i.e. the probability that $y_{i}=-1$.
@@
Appendix B Proofs from Section˜5: hardness of testing E(3,c)(3,c)LIN-2
In this section, we will present a sketch of the proof of the lower bound of testing E(3,c)(3,c)LIN-2. Our main result is stated as follows.
@@ Appendix B Proofs from Section˜5: hardness of testing E(3,c)(3,c)LIN-2
Let us consider the matrix $A\in\{0,1\}^{cN\times N}$ as mentioned in Lemma˜5.4. Based on the matrix $A$, the hard-to-distinguish distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ are as follows:
@@
Note that the system of linear equations in $D_{\mathrm{yes}}$ is satisfiable by setting $x=z$. On the other hand, for the system of linear equations corresponding to $D_{\mathrm{no}}$, vector $y$ is uniformly random. Thus, with high probability, vector AzyAz-y has large Hamming weight for any $z\in\{0,1\}^{N}$, and therefore the system of linear equations $Ax=y$ is far from being satisfiable. The formal proof is in [bogdanov2002lower, Lemma 18]. ∎